OpenStudy (jacobciezki):

Shalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalika will pay? A. B. C. D.

OpenStudy (anonymous):

k back

OpenStudy (jacobciezki):

9/120=x/100 120/x=100/9 x/120=100/9 120/9=100/x

OpenStudy (jacobciezki):

i dont get it

OpenStudy (anonymous):

@iGreen can u help

OpenStudy (anonymous):

jake ill brb agian

OpenStudy (jacobciezki):

ok

OpenStudy (jacobciezki):

hello